How do you solve the problem:
An object is thrown downward with a speed of 16ft/sec from a 160 ft high. The distance it falls in t seconds is s=16t+16t2 (squared). In about how many seconds will it hit the ground?
A. 1.5
B. 2.5
C. 3.5
D. 8.5
Help! Thank you

Answer by jorel1380(2518)

16t2+16t=160
t2+t=10
t(t+1)=10
t equals approximately 2.7 seconds..